OpenStudy (anonymous):

How many gallons of a 80% antifreeze solution must be mixed with 100 gallons of 25% antifreeze to get a mixture that is 70% antifreeze?

OpenStudy (anonymous):

do you know how to start

OpenStudy (anonymous):

No but I know it's a 6 step method

OpenStudy (jdoe0001):

well, we know the amount of 80% of antifreeze will be .. say "x" the amount of 25% of antifreeze is 100 gallons and once we mixed them up, that is "x + 100" they should give use a mix that is 70% antifreeze so \(\begin{array}{llll} &quantity&percentage&total\ with\ antifreeze \\\hline\\ 80\%\ solution&x&0.8&0.8\cdot {\color{brown}{ x}}\\ 25\%\ solution&100&0.25&100\cdot 0.25\\ total\ mixture&{\color{brown}{ x}}+100&0.7&({\color{brown}{ x}}+100)\cdot 0.7 \end{array}\) so... ahemm, what do you think then?

OpenStudy (jdoe0001):

anyhow \(\begin{array}{llll} &quantity&percentage&total\ with\ antifreeze \\\hline\\ 80\%\ solution&{\color{brown}{ x}}&0.8&0.8\cdot {\color{brown}{ x}}\\ 25\%\ solution&100&0.25&100\cdot 0.25\\ total\ mixture&{\color{brown}{ x}}+100&0.7&({\color{brown}{ x}}+100)\cdot 0.7 \end{array}\) so... there, do a system of equations, solve by substitution :)

OpenStudy (jdoe0001):

the SUM of the "total with antifreeze" of the 2 solutions, should equate, the mixture's